OpenStudy (richyw):

Can someone explain to me how to use the Integrating Factor? I have\[u_t=ku_{xx}+Q(x,t)\]on \(-\infty

OpenStudy (richyw):

so I take the FT and get\[\bar{U}_t=-kw^2\bar{U}+\bar{Q}\]and now my solutions say that the "usual" integrating factor can be used to obtain a particular solution. What is this "integrating factor"?

OpenStudy (accessdenied):

Does this look familiar from ordinary DE? \( y' + p(x) y = q(x) \) Integrating Factor: \( \mu = e^{\int p(x) \ dx} \) which has property \(\mu' = p(x) e^{\int p(x) dx} = p(x) \mu(x)\) \( \mu y' + \mu p y = \mu q \) \( \mu y' + \mu' y = \mu q \) \( \left( \mu y \right) ' = \mu q \)

OpenStudy (accessdenied):

Essentially the integrating factor is created with that specific property in mind so that after multiplying both sides by it, the side with y and its first derivative can be pulled together by product rule and then the integral of both sides will remove the combined result.

OpenStudy (richyw):

hmm. i'm not really getting anywhere. my ODE textbook has an explanation of this but i'm getting lost trying to follow the steps.

OpenStudy (richyw):

I have \[\frac{dU}{dt}+kwU=Q(w)\]so I got an integrating factor\[\mu=e^{kw^2/2}\]

OpenStudy (richyw):

then my book tells me to multiply both sides of the equation by this factor.

OpenStudy (richyw):

\[e^{kw^2/2}\frac{dU}{dt}+kwe^{kw^2/2}U=Qe^{kw^2/2}\]

OpenStudy (accessdenied):

You want to look at this equation more like an ordinary differential equation in t. So our integrating factor would also be integrated with respect to t. \( \bar{U} _t = -k \omega ^2\bar{U} + \bar{Q} \) \( \bar{U}_t + k \omega ^2 \bar{U} = \bar {Q} \) We have dU/dt, so we want to multiply everything by a function mainly of t. \( \large \mu (t) = e^{ \int k \omega ^2 dt } \) Note that this looks a bit more like what the end result's exponent has.

OpenStudy (accessdenied):

And the property that I mentioned earlier is that: \( \mu ' (t) = \dfrac{d}{dt} e^{ k \omega ^2 t } = k \omega ^2 e^{k \omega^2 t} = k \omega^2 \color{green}{\mu}\)

OpenStudy (accessdenied):

So multiplying both sides by \( \mu \) yields: \( \mu \bar{U}_t + \color{red}{k \omega ^2 \mu} \bar{U} = \bar{Q} \) This is where the usefulness of integrating factor comes into play.

OpenStudy (accessdenied):

the right side is \( \bar{Q} \mu \) ***

OpenStudy (richyw):

ah, thanks. I think I get it now.

OpenStudy (richyw):

so I end up with like\[\frac{d}{dt}(e^{kw^2t})=Q(w,t)e^{kw^2t}\]

OpenStudy (richyw):

oops \[\frac{d}{dt}(e^{kw^2t}U)=Q(w,t)e^{kw^2t}\]

OpenStudy (accessdenied):

yup, that looks good.

OpenStudy (richyw):

\[e^{kw^2t}U=\int Q(w,t)e^{kw^2t}dt\] \[U=e^{-kw^2t}\int Q(w,t)e^{kw^2t}dt\]

OpenStudy (richyw):

and then the tau is just a dummy variable because the book used a definite integral?

OpenStudy (accessdenied):

Yes. In fact, the definite integral should give you the "particular" solution rather than an indefinite integral with any constants of integration.

OpenStudy (richyw):

right. ok thanks a lot for your help. hopefully my ODE skills don't screw me like this on my final exam!

OpenStudy (accessdenied):

mm, glad to help! and imo integrating factors are the trickiest thing i can remember for methods in ODE. lol